OpenStudy (anonymous):

Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (6, 2) (5, 5) @ma Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (6, 2) (5, 5) @ma @Mathematics

OpenStudy (anonymous):

Solve the following for y:\[\frac{y-2}{x-6}= \frac{y-5}{x-5}\]y=-3x+20

hero (hero):

Oh, that's cool